# SpikeDet: Better Firing Patterns for Accurate and Energy-Efficient Object Detection with Spiking Neuron Networks

### v3

@@ Abstract
Spiking Neural Networks (SNNs) are the third generation of neural networks. They have gained widespread attention in object detection due to their low power consumption and biological interpretability. However, existing SNN-based object detection methods suffer from local firing saturation, where neurons in information-concentrated regions fire continuously throughout all time steps. This abnormal neuron firing pattern reduces the feature discrimination capability and detection accuracy, while also increasing the firing rates that prevent SNNs from achieving their potential energy efficiency. To address this problem, we propose SpikeDet, a novel spiking object detector that optimizes firing patterns for accurate and energy-efficient detection. Specifically, we design a spiking backbone network, MDSNet, which effectively adjusts the membrane synaptic input distribution at each layer, achieving better neuron firing patterns during spiking feature extraction. Additionally, to better utilize and preserve these high-quality backbone features, we introduce the Spiking Multi-direction Fusion Module (SMFM), which realizes multi-direction fusion of spiking features, enhancing the multi-scale detection capability of the model. Experimental results demonstrate that SpikeDet achieves superior performance. On the COCO 2017 dataset, it achieves 51.4% AP, outperforming previous SNN-based methods by 2.5% AP while requiring only half the power consumption. On object detection sub-tasks, including the GEN1 event-based dataset and the URPC 2019 underwater dataset, SpikeDet also achieves the best performance. Notably, on GEN1, our method achieves 47.6% AP, outperforming previous SNN-based methods by 7.2% AP with better energy efficiency.
Index Terms:
Spiking Neuron Network, Object Detection, Neuromorphic Computing.
@@ I Introduction
Object detection is a crucial task in computer vision. It is not only a classification task but also involves regression, requiring the simultaneous identification of object categories and the precise prediction of their locations and sizes [1]. With the development of powerful object detectors [2, 3, 4], this technology has been applied to wide-ranging fields from autonomous driving [5] to intelligent surveillance [6] and medical imaging [7]. However, high-performance detectors require substantial computational resources, while energy-efficient alternatives sacrifice accuracy, posing challenges for resource-constrained deployments [8, 9].
Spiking Neural Networks (SNNs) offer a promising solution to the performance-efficiency trade-off in object detection. Unlike traditional Artificial Neural Networks (ANNs) that rely on continuous values, SNNs communicate between neurons through discrete spiking signals [10]. These spikes, through their temporal dynamic properties, provide SNNs unique information encoding capabilities [11]. Furthermore, this spike-driven mechanism can avoid the computational burden of multiplication operations, instead relying on simpler addition operations. This provides them with significant energy efficiency advantages when deployed on neuromorphic chips, as demonstrated by their implementation on the Speck chip, where power consumption can be as low as 0.7 mW in typical visual processing scenarios [12]. Currently, SNN-based object detection methods have been extensively explored [13, 14, 15, 16, 17, 18, 19, 20, 21], but their performance remains below that of ANN-based object detectors. Moreover, their energy-efficient advantages gradually diminish as their performance improves [20, 21].
@@
Figure 1: Visualization of local firing saturation problem in SNN-based object detector on COCO dataset. Each pixel represents the neuron firing rate, where yellow indicates higher firing rates and blue indicates lower firing rates. (a) and (b) show feature maps before the detection head, as these features determine both classification and regression. (a) averages the $4D$ spike tensor $([T,C,H,W])$ across both time and channel dimensions to show overall spatial firing distribution, while (b) selects a representative channel and averages across time steps to reveal individual neuron firing patterns. (c) shows neuron firing patterns of backbone features at $1/16$ resolution. (d) presents the final detection results.
@@
These limitations motivate us to identify a fundamental problem in existing SNN-based detectors. As shown in Fig. 1, we observe that spiking neurons within bounding boxes are highly active. Particularly in central regions of these boxes, where object information is most concentrated [22, 23], spiking neurons consistently reach maximum firing rates (firing saturation). We define this phenomenon, where multiple adjacent neurons simultaneously exhibit saturated firing patterns, as local firing saturation. These saturated neurons are detrimental for SNN-based object detection, as in SNNs, features are encoded through neuron firing patterns, which fundamentally differ from ANNs. When multiple neurons in these critical areas reach firing saturation, they produce identical representations, significantly reducing feature discriminability capacity. This issue affects both components of object detection, consequently impacting detection accuracy. Specifically, for classification, reduced feature discriminability causes multiple anchor points to produce similar confidence scores, preventing concentrated object representation and degrading overall confidence. For regression, this generates redundant bounding boxes that waste anchor resources and cause missed detections in overlapping scenes, such as the rear person and bench objects in Fig. 1 (d). Additionally, this local firing saturation issue increases the firing rate of SNN detectors, consequently raising their power consumption.
@@
Figure 2: Comparisons with other state-of-the-art methods on COCO AP and Power consumption. Squares represent ANN-based object detectors, while circles represent SNN-based object detectors. (a) Comparison results on the COCO 2017 dataset. (b) Comparison results on the GEN1 dataset.
@@
To address this issue, we propose a novel spiking object detector, SpikeDet, which optimizes the firing patterns of neurons in the model, reducing the probability of local firing saturation to achieve accurate and energy-efficient object detection. The core innovation of our approach is the introduction of a spiking backbone network, Membrane-based Deformed Shortcut Residual Network (MDSNet). Current high-accuracy SNN object detectors [24, 15, 20, 21] employ membrane-based shortcuts [25] to achieve residual learning in their backbone networks for enhancing feature extraction capacity. However, this results in the output of the residual path continuously accumulating directly to the shortcut path. Assuming these outputs are approximately independent, membrane synaptic input variance accumulates across successive layers. This variance amplification significantly increases the probability of neurons receiving extremely large input values. This results in increased frequency of neuron firing saturation, consequently causing the aforementioned local firing saturation problem in information-concentrated regions, as shown in Fig. 1(c). For this, we propose the Membrane-based Deformed Shortcut (MDS), which enables the identity mapping to stabilize its output distribution, fundamentally mitigating the local firing saturation problem. Based on MDS and residual architectures, we design MDSNet, which effectively combines the advantages of both, delivering powerful feature extraction capacity for SNN-based object detectors.
Furthermore, we propose the Spiking Multi-direction Fusion Module (SMFM) to enhance the model’s multi-scale representation capability and preserve the neuron firing patterns of backbone features. In traditional spiking fusion methods [17, 20], information flows along limited paths. Consequently, in overlapping multi-scale object scenarios, shallow layers receive large-scale features from deep layers without adequate processing, causing them to fire saturated across entire regions. Multi-direction fusion creates multiple pathways, allowing features to undergo repeated refinement. This reduces excessive synaptic inputs, mitigating local firing saturation while enhancing multi-scale feature utilization.
@@
(3) Our method achieves better performance. On the COCO 2017 [26], event-based GEN1 [27], and underwater URPC 2019 datasets [28], SpikeDet achieves 51.4%, 47.6%, and 50.1% AP respectively, outperforming other SNN-based approaches while maintaining the lowest power consumption.
The rest of this paper is organized as follows. In Section II, we report the related work. Then we provide the preliminaries of SNNs in Section III. Section IV introduces our proposed method. We present experiment results in Section V. After that, we discuss the hardware deployment of SpikeDet in Section VI. Finally, we conclude this work in Section VII.
@@ II-A Spiking Neural Networks
SNNs are designed to mimic biological behavior more accurately than ANNs through spiking neurons [29]. However, the non-differentiable nature of neuron spike firing prevents SNNs from being trained with traditional backpropagation algorithms used in ANNs. To address this challenge, researchers have proposed two training approaches for SNNs: ANN2SNN conversion and direct training. The ANN2SNN conversion approach approximates ReLU activation using average firing rate, allowing trained ANNs to be transformed into SNNs [30]. Although this approach has led to powerful SNNs [13], it requires thousands of time steps. Moreover, this method is suited only for static datasets, and the performance of converted SNNs depends on the original ANNs. Conversely, the direct training method leverages surrogate gradients to optimize SNNs [31], facilitating their training on a variety of datasets and achieving strong performance within a limited number of time steps. This technique enables the widespread application of SNNs in vision tasks [32, 14, 29]. Hence, we employ the direct training strategy in this study.
To advance the performance of SNNs and realize deeper architectures, researchers have proposed numerous SNN-based residual learning methods [33, 34, 25]. Among them, SEW-ResNet [34] and MS-ResNet [25] address gradient issues and train networks exceeding 100 layers. However, the former does not consider the non-spiking convolutions caused by spike addition, while the latter, despite solving this issue, introduces a local firing saturation problem that we identify in this work. This paper further investigates SNN characteristics to enhance feature extraction while maintaining a full-spiking network.
@@ II-B Object Detection
Object detection is an important and highly complex task in computer vision. Early object detection methods relied on hand-crafted features and sliding window approaches, achieving good accuracy in face detection [35]. With the development of Artificial Neural Networks (ANNs), ANN-based methods gradually became mainstream. These can be categorized into two-stage and one-stage object detectors. Two-stage methods consist of a proposal generator and a region-wise prediction subnetwork [36, 37]. Although they can achieve good performance, they have gradually been replaced by one-stage methods due to their complex design and inefficient inference. One-stage methods directly generate object locations and classifications, significantly reducing computational complexity while maintaining competitive accuracy. Representative one-stage detectors include YOLO [38, 39, 40] and DETR [41, 42] series, which have demonstrated competitive accuracy while maintaining high computational efficiency.
II-C Object Detection with Spiking Neuron Networks
Early attempts to apply SNNs to object detection using ANN2SNN conversion [13, 43] face the challenges of high latency, poor performance, and incompatibility with event cameras. VC-DenseNet [14] and EMS-YOLO [15] pioneer direct training for SNN-based object detection, achieving promising results on both static and event-based datasets. SFOD [17] first introduces an efficient Spiking Fusion Module for SNNs. SpikeYOLO [20] proposes the I-LIF neuron to address the impact of quantization errors from spiking neurons on object detection, significantly improving the performance of SNN-based detectors. However, both approaches overlook the fundamental differences between SNN-based and ANN-based feature representations, i.e., SNN features are represented through neuron firing patterns, which amplifies their performance gap with ANN-based detectors. Hence, our research enhances the SNN object detector by improving firing patterns.
@@ III-A Spiking Neurons
To achieve a balance between biological plausibility and computation complexity in SNNs, researchers have introduced various spiking neural models, including the Hodgkin-Huxley [44], Izhikevich [45], Leaky Integrate-and-Fire (LIF) [46], and Integer Leaky Integrate-and-Fire (I-LIF) [20] models, among others. In our work, we employ the I-LIF neuron due to its effectiveness in alleviating the quantization error problem in SNNs. The I-LIF neuron model is represented by:
$\mathbf{u}^{t,n}=\tau\mathbf{h}^{t-1,n}+\mathbf{x}^{t,n-1},$ (1)
[…]
@@
[…]
Proposition 1. The integer spikes output by I-LIF neurons during training can be converted to the sum of binary spikes from SR-IF neurons over D time steps during inference, provided that these SR-IF neurons are configured with: (1) a threshold $V_{th}=1$, and (2) their inputs $\mathbf{x}^{t,n-1}$ adjusted by a bias of $0.5(1-\tau_{I-LIF})^{t-1}$.
Proof. In SR-IF neurons, the differences from I-LIF neurons lie in their integration and firing mechanisms. Specifically, in the integration process, the decay constant $\tau_{SR-IF}=1$, and their firing mechanism changes to
$\mathbf{o}^{t,n}=Hea\left(\mathbf{u}^{t,n}-V_{th}\right),$ (4)
where $Hea$ represents the Heaviside step function. When converting I-LIF neurons to SR-IF neurons over D time steps, the inputs must be consistent. We can set the input to the SR-IF neuron at time step $t=1$ as the original input, while maintaining zero inputs for time steps $t=2,...,D$. Then, with $V_{th}=1$, we can derive,
$\mathbf{o}^{t,n}_{I-LIF}$ $=\sum_{d=1}^{D}Hea\left(\mathbf{u}^{d,n}_{SR-IF}-1\right)$ (5)
$=floor(Clip(\mathbf{u}^{1,n}_{SR-IF},0,D))$
$=Clip(floor(\mathbf{u}^{1,n}_{SR-IF}),0,D)$
$=Clip(round(\mathbf{u}^{1,n}_{SR-IF}-0.5),0,D).$
To make Eq.5 equivalent to Eq.2, we need to ensure that $\mathbf{u}^{t,n}_{I-LIF}=\mathbf{u}^{1,n}_{SR-IF}-0.5$. To achieve this, we must add appropriate bias to the synaptic inputs of the neuron. According to Eqs.1 and Eq.3, we need to add a bias of $0.5(1-\tau_{I-LIF})^{t-1}$ to $\mathbf{x}^{t,n-1}$.
Therefore, by performing this conversion during inference, it maintains the spike-driven computation mechanism while reducing quantization errors. Furthermore, to ensure that $\tau_{I-LIF}$ can function effectively, we employ Soft-Reset LIF neurons at the time steps $[1,...,(t-1)\times D,t\times D,...,T\times D]$ when it is deployed. This does not affect the conclusion of Proposition 1, as it only influences how temporal information from the previous time step $t$ connects to the current one, without impacting the integer spike conversion process.
@@
Figure 3: The architecture of SpikeDet. SpikeDet comprises MDSNet, SMFM, and the SpikeYOLO Detection Head [20]. The model receives two types of inputs: event and static data, with the input coding and output represented in the figure. The core of SpikeDet is MDSNet, which consists of 5 stages. Each stage downsampling factor relative to the original input is marked in the figure, with MDS-Block1 and MDS-Block2 performing feature extraction without and with downsampling, respectively. By incorporating MDS, this architecture successfully stabilizes the firing patterns at each stage, alleviating the local firing saturation problem. Additionally, SMFM enables multi-direction feature fusion, allowing features to undergo multiple refinements through the model, thereby improving the model’s capability to detect multi-scale objects and preserving neuron firing patterns.
III-B Power Consumption
The energy efficiency of SNNs primarily comes from performing Accumulation Calculations (ACs) only when neurons fire. However, many current SNN-based works cannot guarantee a full-spiking network. Therefore, when calculating the energy consumption of these networks, we also consider Multiplication and Addition Calculations (MACs). For ANNs, as the vast majority of their operations are MACs with only a small number of ACs, we ignore these ACs operations in our calculations. Furthermore, in line with previous works [17, 15, 47, 19, 21], we quantify the specific power consumption values using $E_{\text{MACs}}=4.6\text{pJ}$ and $E_{\text{ACs}}=0.9\text{pJ}$. The power consumption calculation formulas for SNNs and ANNs are shown respectively as follows, where $fr$ represents the firing rate of the model, $T$ denotes the number of time steps, and $N$ is the number of operations.
$E_{SNNs}=T\times(fr\times E_{ACs}\times N_{ACs}+E_{MACs}\times N_{MACs})$ (6)
$E_{ANNs}=T\times E_{MACs}\times N_{MACs}$ (7)

### v4

# SpikeDet: Better Firing Patterns for Accurate and Energy-Efficient Object Detection with Spiking Neural Networks

@@ Abstract
Spiking Neural Networks (SNNs) are the third generation of neural networks. They have gained widespread attention in object detection due to their low power consumption and biological interpretability. However, existing SNN-based object detection methods suffer from local firing saturation, where adjacent neurons concurrently reach maximum firing rates, especially in object-centric regions. This abnormal neuron firing pattern reduces the feature discrimination capability and detection accuracy, while also increasing the firing rates that prevent SNNs from achieving their potential energy efficiency. To address this problem, we propose SpikeDet, a novel spiking object detector that optimizes firing patterns for accurate and energy-efficient detection. Specifically, we design a spiking backbone network, MDSNet, which effectively adjusts the membrane synaptic input distribution at each layer, achieving better neuron firing patterns during spiking feature extraction. For the neck, to better utilize and preserve these high-quality backbone features, we introduce the Spiking Multi-direction Fusion Module (SMFM), which realizes multi-direction fusion of spiking features, enhancing the multi-scale detection capability of the model. Furthermore, we propose the Local Firing Saturation Index (LFSI) to quantitatively measure local firing saturation. Experimental results validate the effectiveness of our method, with SpikeDet achieving superior performance. On the COCO 2017 dataset, it achieves 52.2% AP, outperforming previous SNN-based methods by 3.3% AP while requiring only half the power consumption. On object detection sub-tasks, including event-based GEN1, underwater URPC 2019, low-light ExDARK, and dense scene CrowdHuman datasets, SpikeDet also achieves the best performance.
@@ I Introduction
Spiking Neural Networks (SNNs) have emerged as a promising alternative to traditional Artificial Neural Networks (ANNs) [huang2025converting]. Unlike ANNs that rely on continuous values, spiking neurons in SNNs emulate biological neurons, computing and communicating through discrete spiking signals [maass_networks_1997]. This spike-driven mechanism can avoid the computational burden of multiplication operations, instead relying on simpler addition operations, providing significant potential for energy-efficient computing [zhang2024accurate]. Furthermore, the temporal dynamics of these spikes provide SNNs with distinct representation capabilities [skatchkovsky_learning_2021]. Motivated by these characteristics, SNNs have been increasingly applied to various computer vision tasks [cao2023event, cordone_object_2022, barchid2023spiking, safa2021improving].
Among these, SNNs have achieved competitive performance with ANNs on relatively simple tasks such as image classification [yao_spikedriven_2024, yao_scaling_2025]. However, for more complex and widely applied tasks like object detection, which require simultaneous identification of object categories and precise prediction of their locations and sizes [liu2019pay, zhou2024dpnet], SNNs still face significant challenges. Specifically, performance gaps persist between SNN-based and ANN-based detectors [kim_spikingyolo_2020, cordone_object_2022, su_deep_2023, yu_spikingvit_2024, fan_sfod_2024, qu_spiking_2024, wang_eassnn_2025, luo_integervalued_2025, yao_scaling_2025]. More critically, the energy efficiency advantages of SNNs gradually diminish as their performance improves [luo_integervalued_2025, yao_scaling_2025].
@@
Figure 1: Visualization of local firing saturation problem in SNN-based object detector on COCO dataset. Each pixel represents the neuron firing rate. (a) and (b) show feature maps before the detection head, as these features determine both classification and regression. (a) averages the $4D$ spike tensor $([T,C,H,W])$ across both time and channel dimensions to show overall spatial firing distribution, while (b) selects a representative channel and averages across time steps to reveal individual neuron firing patterns. (c) shows neuron firing patterns of backbone features at $1/16$ resolution. (d) presents the final detection results.
@@
These limitations motivate us to identify a fundamental problem in existing SNN-based detectors. As shown in Fig. 1, we observe that spiking neurons consistently reach maximum firing rates (firing saturation) in certain spatial regions, particularly in central regions of bounding boxes where object information is most concentrated [tian_fcos_2019, zhang_bridging_2020]. We define this phenomenon, where multiple spatially adjacent neurons concurrently exhibit saturated firing patterns, as local firing saturation. These saturated neurons are detrimental for SNN-based object detection. In SNNs, features are encoded through neuron firing patterns. When multiple adjacent neurons reach firing saturation, they produce identical representations, significantly reducing feature discriminability. This issue affects both components of object detection, consequently impacting detection accuracy. Specifically, for classification, reduced feature discriminability causes multiple anchor points to produce similar confidence scores, preventing concentrated object representation and degrading overall confidence. For regression, this generates redundant bounding boxes that waste anchor resources and cause missed detections in overlapping scenes, such as the rear person and bench objects in Fig. 1 (d). Additionally, this local firing saturation issue increases the firing rate of SNN detectors, consequently raising their power consumption.
@@
Figure 2: Comparisons with other state-of-the-art methods on COCO AP and Power consumption. Squares represent ANN-based object detectors, circles represent SNN-based object detectors, and triangles represent our methods. (a) Comparison results on the COCO 2017 dataset. (b) Comparison results on the GEN1 dataset.
@@
To address this issue, we propose SpikeDet, a novel spiking object detector. By optimizing neuron firing patterns, SpikeDet mitigates local firing saturation and achieves accurate, energy-efficient object detection. The core innovation of our approach is the introduction of a spiking backbone network, Membrane-based Deformed Shortcut Residual Network (MDSNet). Current high-accuracy SNN object detectors [yao_spikedriven_2024, su_deep_2023, luo_integervalued_2025, yao_scaling_2025] employ membrane-based shortcuts [hu_advancing_2024] to achieve residual learning in their backbone networks for enhancing feature extraction capacity. However, this results in the residual path output being directly added to the shortcut path at each layer. As we theoretically analyze in Section IV-B1, such repeated path summation causes membrane synaptic input variance to accumulate across successive layers. This variance amplification significantly increases the probability of neurons receiving excessively large synaptic inputs. Consequently, neurons more frequently reach firing saturation. Combined with the inherently high neural activity in information-concentrated regions, this causes adjacent neurons to concurrently saturate, leading to the local firing saturation problem shown in Fig. 1(c). For this, we propose the Membrane-based Deformed Shortcut (MDS), which enables the identity mapping to stabilize its output distribution, fundamentally mitigating the local firing saturation problem. Based on MDS and residual architectures, we design MDSNet, which effectively combines the advantages of both, delivering powerful feature extraction capacity for SNN-based object detectors.
Multi-scale feature fusion also plays an important role in mitigating local firing saturation. For this, we propose the Spiking Multi-direction Fusion Module (SMFM). In traditional spiking fusion methods [fan_sfod_2024, luo_integervalued_2025], information flows along limited paths. Such insufficient fusion raises two issues. First, features from different scales may simultaneously activate at object centers. Limited fusion prevents their effective integration, causing their responses to accumulate and leading to local firing saturation. Second, with limited fusion, shallow detail-sensitive features lack sufficient semantic modulation when integrated with deeper features, causing local firing saturation at regions with high local contrast in the fused representations. Multi-direction fusion creates multiple pathways, allowing features to undergo repeated refinement. This allows better integration of information across scales, mitigating local firing saturation while enhancing multi-scale feature utilization. Additionally, to quantitatively evaluate the capability of our method in addressing local firing saturation, we propose the Local Firing Saturation Index (LFSI). This metric averages the proportion of saturated neurons surrounding each neuron across all layers, reflecting the severity of local firing saturation in the network.
@@
(3) Our method achieves superior performance, and its effectiveness in addressing local firing saturation is quantitatively verified by the proposed LFSI. On the COCO 2017 dataset [lin_microsoft_2014] and various object detection sub-tasks, SpikeDet outperforms other SNN-based approaches while maintaining the lowest power consumption.
The rest of this paper is organized as follows. In Section II, we report the related work. Then we provide the preliminaries of SNNs in Section III. Section IV introduces our proposed method. We present experiment results in Section V. Finally, we conclude this work in Section VI.
@@ II-A Spiking Neural Networks
SNNs are designed to mimic biological behavior more accurately than ANNs through spiking neurons [yao_spikedriven_2024]. However, the non-differentiable nature of neuron spike firing prevents SNNs from being trained with traditional backpropagation algorithms used in ANNs. To address this challenge, researchers have proposed two training approaches for SNNs: ANN2SNN conversion and direct training. The ANN2SNN conversion approach approximates ReLU activation using average firing rate, allowing trained ANNs to be transformed into SNNs [huang2025converting]. Although this approach has led to powerful SNNs [kim_spikingyolo_2020], it requires thousands of time steps. Moreover, this method is suited only for static datasets, and the performance of converted SNNs depends on the original ANNs. Conversely, the direct training method leverages surrogate gradients to optimize SNNs [fan_sfod_2024], facilitating their training on a variety of datasets and achieving strong performance within a limited number of time steps. Therefore, we employ the direct training strategy in this study.
To advance the performance of SNNs and realize deeper architectures, researchers have proposed numerous SNN-based residual learning methods [zheng_going_2021, fang_deep_2021, hu_advancing_2024]. Among them, SEW-ResNet [fang_deep_2021] and MS-ResNet [hu_advancing_2024] address gradient issues and train networks exceeding 100 layers. However, the former does not consider the non-spiking convolutions caused by spike addition. The latter solves this issue but introduces a local firing saturation problem that we identify in this work. This paper further investigates SNN characteristics to enhance feature extraction while maintaining a full-spiking network.
@@ II-B Object Detection
Object detection is an important and highly complex task in computer vision. Early object detection methods relied on hand-crafted features and sliding window approaches, achieving good accuracy in face detection [viola_rapid_2001]. With the development of Artificial Neural Networks (ANNs) [yin2025semi, yin2025uncertainty], ANN-based methods gradually became mainstream. These can be categorized into two-stage and one-stage object detectors. Two-stage methods consist of a proposal generator and a region-wise prediction subnetwork [li2023proposal]. Although they can achieve good performance, they have gradually been replaced by one-stage methods due to their complex design and inefficient inference. One-stage methods directly generate object locations and classifications, significantly reducing computational complexity while maintaining competitive accuracy. Representative examples include YOLO [yolov8_ultralytics, wang_goldyolo_2023, yolo11_ultralytics] and DETR [carion_endtoend_2020] series.
II-C Object Detection with Spiking Neural Networks
Early attempts to apply SNNs to object detection using ANN2SNN conversion [kim_spikingyolo_2020, kim_fast_2020] face the challenges of high latency, poor performance, and incompatibility with event cameras. VC-DenseNet [cordone_object_2022] and EMS-YOLO [su_deep_2023] pioneer direct training for SNN-based object detection, achieving promising results on both static and event-based datasets. SFOD [fan_sfod_2024] first introduces an efficient Spiking Fusion Module for SNNs. SpikeYOLO [luo_integervalued_2025] proposes the I-LIF neuron to address the impact of quantization errors from spiking neurons on object detection, significantly improving the performance of SNN-based detectors. However, both approaches overlook the fundamental differences between SNN-based and ANN-based feature representations, i.e., SNN features are represented through neuron firing patterns. This oversight amplifies their performance gap with ANN-based detectors. Hence, our research enhances the SNN object detector by improving neuron firing patterns.
@@ III-A Spiking Neurons
To achieve a balance between biological plausibility and computation complexity in SNNs, researchers have introduced various spiking neural models, including the Hodgkin-Huxley [hodgkin_quantitative_1952], Integrate-and-Fire (IF) [gerstner2014neuronal], Leaky Integrate-and-Fire (LIF) [abbott_lapicques_1999a], and Integer Leaky Integrate-and-Fire (I-LIF) [luo_integervalued_2025] models, among others. In our work, we employ the I-LIF neuron due to its effectiveness in reducing the quantization error problem of SNNs. The I-LIF neuron model is represented by:
$\mathbf{u}^{t,n}=\tau\mathbf{h}^{t-1,n}+\mathbf{x}^{t,n},$ (1)
$\mathbf{o}^{t,n}=\mathit{Clip}\left(\mathit{round}(\mathbf{u}^{t,n}),0,D\right),$ (2)
@@
Here, $\mathbf{u}$ represents the membrane potential, $\mathbf{h}$ the membrane potential retained from the previous time step, $\mathbf{x}$ the pre-synaptic input, and $\mathbf{o}$ the output spike. Variables $t$ and $n$ denote time step and layer number, respectively. Function $\mathit{Clip}\left(x,min,max\right)$ constrains $x$ within $[min,max]$, while $\mathit{round}(x)$ rounds $x$ to the nearest integer. $D$ is the upper limit hyperparameter for integer-valued spike activation, $\tau$ is the membrane potential decay constant, and $V_{th}$ is the membrane potential threshold. The core of the I-LIF neuron lies in Eq. 2, which enables neurons to generate integer spikes during training and can convert them to the sum of binary spikes in $D$ time steps of Soft-Reset IF neurons during inference [yao_scaling_2025].
@@
Figure 3: The architecture of SpikeDet. SpikeDet comprises MDSNet, SMFM, and the SpikeYOLO Detection Head [luo_integervalued_2025]. The model receives two types of inputs: event and static data, with the input coding and output represented in the figure. The core of SpikeDet is MDSNet, which consists of 5 stages. Each stage downsampling factor relative to the original input is marked in the figure, with MDS-Block1 and MDS-Block2 performing feature extraction without and with downsampling, respectively. By incorporating MDS, this architecture successfully stabilizes the firing patterns at each stage, alleviating the local firing saturation problem. Additionally, SMFM enables multi-direction feature fusion, allowing features to undergo multiple refinements through the model, thereby improving the model’s capability to detect multi-scale objects and preserving neuron firing patterns.
III-B Surrogate Gradient Training Strategy
In SNN training, a key challenge arises from the non-differentiable nature of the spike generation process described in Eq. 2. To overcome this limitation and enable effective backpropagation within contemporary deep learning frameworks, we implement surrogate gradient techniques. The surrogate gradient applied to Eq. 2 is expressed as:
$\frac{\partial\mathbf{s}^{t,n}}{\partial\mathbf{u}^{t,n}}=\begin{cases}1,&\text{if }0\leq\mathbf{u}^{t,n}\leq D\\ 0,&\text{otherwise.}\end{cases}$ (4)
For common LIF neurons, the spike firing formula is
$\mathbf{o}^{t,n}=Hea\left(\mathbf{u}^{t,n}-V_{th}\right).$ (5)
The corresponding surrogate gradient is expressed as:
$\frac{\partial\mathbf{o}^{t,n}}{\partial\mathbf{u}^{t,n}}=\frac{1}{a}\text{sign}\left(\left|\mathbf{u}^{t,n}-V_{\text{th}}\right|\leq\frac{a}{2}\right),$ (6)
with $a$ serving as a normalization coefficient that guarantees the function integrates to 1.